OpenStudy (anonymous):

Given two component vectors, solve for the resultant vector. Use pythagorom theorm to find the hypotenuse and use inverse (arc) tangent to solve for the angle 1. x=-.75 y=-1.25 2. x=20000 y=14000

OpenStudy (anonymous):

\[\left|\bar{r}\right|=\sqrt{x^2+y^2}\]for the magnitude of the resultant vector\[\theta=\tan^{-1}\left(\frac{y}{x}\right)\]for the angle with reference to x-axis

OpenStudy (anonymous):

... the sign of x and y indicates the quadrant the vector exist\[\bar{r}=\left|\bar{r}\right|\angle\theta\]the resultant vector

OpenStudy (anonymous):

so what did you get as your answers?